# FedMLAC: Mutual Learning Driven Heterogeneous Federated Audio Classification

### v1

@@ Abstract
Federated Learning (FL) provides a privacy-preserving paradigm for training audio classification (AC) models across distributed clients without sharing raw data. However, Federated Audio Classification (FedAC) faces three critical challenges that substantially hinder performance: data heterogeneity, model heterogeneity, and data poisoning. While prior works have attempted to address these issues, they are typically treated independently, lacking a unified and robust solution suited to real-world federated audio scenarios. To bridge this gap, we propose FedMLAC, a unified mutual learning framework designed to simultaneously tackle these challenges in FedAC. Specifically, FedMLAC introduces a dual-model architecture on each client, comprising a personalized local AC model and a lightweight, globally shared Plug-in model. Through bidirectional knowledge distillation, the Plug-in model enables global knowledge transfer while adapting to client-specific data distributions, thus supporting both generalization and personalization. To further enhance robustness against corrupted audio data, we develop a Layer-wise Pruning Aggregation (LPA) strategy that filters unreliable Plug-in model updates based on parameter deviations during server-side aggregation. Extensive experiments on four diverse audio classification benchmarks—spanning both speech and non-speech tasks—demonstrate that FedMLAC consistently outperforms existing state-of-the-art methods in terms of classification accuracy and robustness to noisy data.
@@ 1 Introduction
Audio classification (AC), defined as the task of assigning semantic labels to audio signals, has become a foundational component of modern intelligent systems [1]. It supports a wide range of applications, including speech command recognition [2], emotion detection [3], environmental sound classification, and bioacoustic monitoring [4]. These tasks span diverse acoustic domains and require models to capture fine-grained temporal and spectral features from highly variable and often noisy data [5, 6]. Advances in deep learning—particularly convolutional neural networks (CNNs), recurrent neural networks (RNNs), and transformer-based architectures—have significantly improved classification performance under both supervised and weakly supervised settings [1], relying on large-scale, diverse datasets to learn robust acoustic representations from raw waveforms or spectrograms. In practice, however, such datasets are increasingly generated and retained locally on heterogeneous edge devices such as smartphones, wearables, smart speakers, and IoT sensors [7], rather than being centrally collected. This shift toward decentralized data generation poses both technical and ethical challenges to centralized training, especially in light of privacy concerns, communication overhead, and legal regulations such as the GDPR [8] and the EU AI Act [9]. These constraints underscore the need for alternative learning frameworks that can operate effectively in decentralized environments while ensuring data privacy and communication efficiency.
One of the most promising solutions to address the challenges of decentralized and privacy-sensitive audio data is Federated Learning (FL), a distributed training paradigm that enables multiple clients to collaboratively learn a shared model without exchanging raw data [10]. By performing computation locally and aggregating only model updates, FL reduces communication overhead and mitigates privacy risks, making it particularly attractive for deployment on edge devices such as smartphones, wearables, and IoT sensors [11, 12]. This paradigm aligns well with the characteristics of modern audio data ecosystems, where large volumes of diverse audio signals remain distributed across heterogeneous sources [6]. Compared to centralized learning approaches, FL offers inherent advantages in terms of privacy preservation, regulatory compliance, and scalability [13]. These properties make it a natural candidate for large-scale AC tasks under real-world constraints.
In this context, Federated Audio Classification (FedAC) refers to the application of FL to collaboratively train AC models across distributed clients while keeping local audio data private [14, 6, 15, 16, 17]. FedAC enables learning from diverse, sensitive, and geographically dispersed audio sources without requiring centralized access, which is particularly valuable in domains such as healthcare [18], smart environments [16], and personalized voice interfaces [19]. However, applying FL to AC introduces several unique challenges. First, data heterogeneity arises from variations in speaker identity, dialect, background noise, and recording conditions, resulting in highly non-IID data distributions across clients [20, 21, 22]. Second, model heterogeneity stems from differences in device capabilities, runtime environments, and model architectures, complicating coordination and degrading global convergence [23, 24]. Third, data poisoning in the federated setting may arise from both malicious clients and naturally occurring issues such as label errors or input noise, potentially leading to degraded global model performance and unstable convergence [25, 26]. These challenges collectively limit the performance, robustness, and scalability of conventional FL approaches when applied to real-world AC tasks.
@@
Despite growing interest in FL for both speech and non-speech audio applications, existing research tends to address the challenges of data heterogeneity, model heterogeneity, and data poisoning in partial or disjointed ways. A substantial body of work focuses on mitigating non-IID data distributions through strategies such as regularization techniques [27, 28], client selection [29, 30, 31, 32, 33, 34], representation learning [35], or robust aggregation mechanisms [36, 20, 21, 37]. While these techniques improve global model utility under skewed data, they often assume a homogeneous model architecture shared across all clients, which is impractical in real-world deployments involving heterogeneous edge devices. To address system-level diversity, recent studies have explored flexible model configurations, including personalized layers [38, 39, 40], decoupled training [41, 27, 42, 43, 16], incremental learning [44, 45, 46], or sparsified aggregation [47, 48]. However, such methods rarely account for the interaction between model and data heterogeneity, which can cause unstable convergence, degraded generalization, or fairness concerns when local models diverge significantly.
@@
Another important challenge in FedAC is the presence of corrupted or unreliable local data, commonly referred to as data poisoning. Such corruption can arise not only from adversarial clients but also from natural sources, including label inconsistencies in subjective tasks (e.g., emotion recognition) and distorted audio recordings in uncontrolled environments. These issues can significantly impair model stability and convergence, especially when coupled with non-IID data or heterogeneous client models. To mitigate these effects, researchers have proposed defense strategies such as anomaly detection [27, 49, 35, 50], adversarial learning [51], client clustering [52], and model pruning [53, 54]. While effective in isolation, these methods are typically reactive and rarely consider the joint effects of data and model heterogeneity. Moreover, many introduce a trade-off between robustness and utility, often degrading performance on clean or low-resource clients. Thus, achieving resilience to data corruption without sacrificing personalization or fairness remains an open problem in FedAC.
@@
[36, 20, 21, 29, 35, 37, 28, 33, 32] $\checkmark$ $\times$ $\times$
[47, 39, 48, 40, 42, 43] $\times$ $\checkmark$ $\times$
@@
[38, 41, 44, 45, 46] $\checkmark$ $\checkmark$ $\times$
@@
As summarized in Table 1, existing research efforts in FedAC remain largely fragmented, with most methods targeting only one or two specific challenges. While notable progress has been made in mitigating data heterogeneity or accommodating model diversity, few approaches recognize their interdependence, and even fewer address the added complexity introduced by corrupted or unreliable local data. This gap is especially consequential in AC scenarios, where task subjectivity, noisy environments, and hardware diversity are the norm rather than the exception. The interplay between statistical non-IID distributions, architectural heterogeneity, and data poisoning presents a compounded threat to model convergence, generalization, and fairness—making partial solutions insufficient for deployment in dynamic, real-world settings. To the best of our knowledge, no existing work proposes a unified framework capable of jointly addressing all three challenges in a coordinated and scalable manner. This observation underlines a critical need for integrated approaches that go beyond siloed techniques and holistically enhance the robustness and adaptability of FedAC systems.
To address the intertwined challenges of data heterogeneity, model heterogeneity, and data corruption in FedAC, we propose FedMLAC, a Mutual Learning-driven Federated Audio Classification framework. FedMLAC consists of two key components: FedMLAC-Update and FedMLAC-Aggregation. In FedMLAC-Update, each client maintains a personalized local AC model and a lightweight server-distributed Plug-in model. Through mutual learning, the Plug-in model transfers global knowledge to the local model while simultaneously absorbing client-specific information. This bidirectional distillation ensures that clients benefit from global generalization and local personalization without sharing raw data. Meanwhile, in FedMLAC-Aggregation, the updated Plug-in models from clients are aggregated on the server via a robust Layer-wise Pruning Aggregation (LPA) strategy. This approach selectively integrates trustworthy model layers while pruning extreme deviations—caused by non-IID data or adversarial noise—thereby improving the resilience and stability of the global Plug-in model. Together, these two modules enable FedMLAC to achieve scalable, personalized, and communication-efficient learning across heterogeneous audio environments.
A detailed comparison between FedMLAC and representative federated approaches is presented in Table 1, highlighting the framework’s ability to simultaneously tackle statistical data heterogeneity, model diversity, noisy or corrupted data, and personalization—four challenges that prior methods typically address in isolation. FedMLAC unifies these components into a cohesive and scalable framework: it supports flexible, personalized client models via decoupled optimization; enhances local adaptation and global generalization through mutual knowledge distillation; and achieves robust model aggregation via layer-wise pruning to filter out anomalous or adversarial updates. This integrated design allows FedMLAC to maintain strong performance across highly heterogeneous and dynamic AC environments, as verified in our extensive experimental evaluation.
The main contributions of this work are summarized as follows:
  • 1.
    We propose FedMLAC, a unified federated learning framework for audio classification that is capable of simultaneously addressing three fundamental challenges: data heterogeneity, model heterogeneity, and data poisoning. To the best of our knowledge, this is the first framework to jointly solve all three challenges in the federated audio classification setting.
  • 2.
    To effectively tackle data and model heterogeneity, we introduce a mutual learning mechanism based on a lightweight, globally shared Plug-in model. This Plug-in model enables bidirectional knowledge transfer between global and local models, enhancing generalization across clients while supporting flexible, personalized architectures.
  • 3.
    To enhance robustness against data poisoning—particularly from data noise and label errors—we propose a Layer-wise Pruning Aggregation (LPA) strategy. LPA identifies and prunes abnormal Plug-in model updates at the layer level by measuring parameter deviations across clients, effectively filtering out unreliable contributions during aggregation. This mechanism improves the stability and resilience of the global Plug-in model under noisy and inconsistent local data conditions.
  • 4.
    We provide a theoretical convergence analysis of FedMLAC, showing that the global Plug-in model converges to a stationary point at a rate of $\mathcal{O}(1/\sqrt{T})$ under standard non-convex assumptions. The analysis rigorously characterizes the impact of data heterogeneity, stochastic variance, and layer-wise pruning deviations, thereby offering theoretical guarantees for FedMLAC’s robustness and scalability in heterogeneous federated environments.
  • 5.
    We conduct extensive experiments on four benchmark audio classification datasets, encompassing both speech and non-speech tasks. The results demonstrate that FedMLAC consistently outperforms existing state-of-the-art methods in terms of classification accuracy and robustness to noisy data.
The remainder of this paper is organized as follows. Section 2 provides a comprehensive review of related work on FedAC. Section 3 introduces the problem formulation. Section 4 details the core components and technical implementation of the proposed approach. Section 6 presents the experimental setup, results, and analyses conducted across multiple benchmark datasets. A formal convergence analysis is offered in Section 5. Finally, Section 7 summarizes the key findings and discusses potential directions for future research.
2 Related Work
2.1 Audio Classification
Audio classification (AC) involves automatically assigning predefined semantic labels to audio recordings, such as speech commands, music genres, or environmental sounds. Traditional techniques largely relied on handcrafted features like MFCCs and classical classifiers (e.g., SVMs, GMMs) [55, 56]. With the rise of deep learning, architectures such as CNNs and RNNs have shown superior performance by learning hierarchical representations from raw or spectrogram-based audio [57]. Convolutional Recurrent Neural Networks (CRNNs), which combine the spatial feature extraction power of CNNs with the temporal modeling capabilities of RNNs, have become a widely adopted architecture in audio tagging and event detection [58]. CRNNs typically consist of several convolutional layers for capturing time-frequency local patterns, followed by recurrent layers (e.g., bidirectional GRUs or LSTMs) to model long-term temporal dependencies, and fully connected output layers for final prediction [59].
Despite these advancements, large-scale, high-quality audio datasets remain difficult to collect due to the need for manual annotation and concerns around user privacy. Speech and ambient sound recordings often contain personally identifiable or context-sensitive content, posing regulatory and ethical challenges under data protection frameworks like GDPR [8]. Furthermore, real-world datasets are frequently noisy or weakly labeled, especially when sourced from web platforms [60]. These issues highlight the need for classification systems that are robust to imperfect data, and capable of learning from decentralized or privacy-sensitive sources—motivating recent interest in distributed machine learning paradigms.
2.2 Federated Learning
Federated learning (FL) is a decentralized machine learning paradigm that enables multiple clients to collaboratively train a global model without sharing raw data, thereby preserving user privacy [10, 13]. Its privacy-preserving nature has led to widespread adoption in sensitive domains such as healthcare, mobile services, and finance [13]. However, FL is fundamentally constrained by statistical heterogeneity, where non-IID client data distributions impair convergence and model generalization [11]. To alleviate this, the literature has explored techniques such as regularization-based local training [61, 62], robust aggregation mechanisms [63, 64], and personalized learning frameworks [65]. Simultaneously, the threat of data poisoning—where adversarial clients compromise model integrity—has motivated Byzantine-resilient strategies [66] and trust-aware aggregation schemes [13]. Despite this progress, FL research remains predominantly focused on vision and text benchmarks like CIFAR-10, FEMNIST, and GLUE, with relatively limited efforts targeting audio-based tasks [13]. Given the unique temporal and spectral complexities of audio signals and the privacy sensitivities inherent in speech data, extending FL to acoustic learning scenarios presents both urgent demand and technical novelty. Our work addresses this gap by proposing a federated audio classification framework that accounts for these domain-specific challenges.
2.3 Federated Audio Classification
Recent advancements in Federated Audio Classification (FedAC) have progressively addressed the inherent challenges of decentralized learning over acoustic data. Early efforts in this field largely adopted techniques from computer vision and speech recognition, applying them directly to audio domains without explicitly tailoring to their temporal, spectral, or noise characteristics. For instance, works such as [27, 28] focused on mitigating data heterogeneity through regularization-based federated training, while [29, 30, 32] explored client selection or clustering to group clients with similar data distributions. Representation learning strategies have also emerged [35] to align feature spaces across clients, showing promise in noisy acoustic environments.
More recently, system heterogeneity—stemming from the diversity of devices, compute resources, and model capacities—has gained attention. Personalized FL schemes like [38, 39] allow clients to retain unique local layers, while parameter decoupling [41, 27] and sparse aggregation [47, 48] seek to reduce communication and computation loads. Incremental training paradigms [45, 46] further enhance long-term adaptability, especially under shifting client distributions.
Simultaneously, growing attention has turned to the issue of label noise and unreliable audio data. Both adversarial poisoning [51] and benign corruption—e.g., in emotion recognition or field-recorded data—can derail global convergence. In response, anomaly detection [49, 50], model pruning [53, 54], and unsupervised clustering [52] have been proposed. However, most of these strategies remain isolated to specific challenges, often neglecting the complex interplay between noisy data, non-IID distributions, and heterogeneous model architectures. Few studies attempt to jointly address all three, which leads to limited generalization, unstable training, or diminished fairness on underrepresented clients. As such, holistic frameworks that simultaneously account for statistical, structural, and adversarial variations remain critically underexplored in the current FedAC landscape.
3 Prelimineries
3.1 Problem Setup
@@
where $n=\sum_{k=1}^{N}{n_{k}}$, $n_{k}=|D_{k}|$, $\mathcal{L}_{k}(\omega^{k})=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell(\mathbf{\omega^{k}};\mathbf{x}_{k,j},\mathbf{y}_{k,j})$ denotes the average loss of the ${k^{th}}$ client model $\omega^{k}$ over $D_{k}$, and $l(\cdot)$ represents the training loss (e.g., cross-entropy (CE) loss for AC tasks).
@@
$\omega^{k}_{t+(i+1)}=\omega^{k}_{t+(i)}-{\eta_{t+(i)}}\nabla{\mathcal{L}_{k}}(\omega^{k}_{t+(i)}),i=0,1,...,E-1,$ (2)
where $\eta$ is the learning rate, $\omega_{t}^{k}$ represents the $k^{th}$ client model in the $t^{th}$ communication round. and $E$ denotes the local training epochs.
After local training, each client sends its updated model $\mathbf{\omega}_{t+1}^{k}$ to the central server, which performs aggregation to update the global model. The most commonly used aggregation method is Federated Averaging (FedAvg) [10], defined as:
@@
The basic FedAC framework requires that each client’s AC model $\mathbf{\omega}^{k}$ maintains the same model architecture. Furthermore, the heterogeneity of distributed AC data leads to an inconsistency between the learning objectives of client models and the global FedAC training objective. These challenges significantly reduce the practicality and stability of existing FedAC frameworks.
3.2 Mutual Learning
@@
$\mathcal{L}_{i}&=\mathcal{L}_{\text{task}}(\mathbf{\omega}^{i};\mathcal{D}_{i})+\lambda\mathcal{L}_{\text{KL}}(\mathit{p}_{i},\mathit{p}_{j})\\ \mathcal{L}_{j}&=\mathcal{L}_{\text{task}}(\mathbf{\omega}^{j};\mathcal{D}_{j})+\lambda\mathcal{L}_{\text{KL}}(\mathit{p}_{j},\mathit{p}_{i})\end{aligned}\right.,$ (1)
@@
where $\mathit{p}_{i}(c)$ represents the probability of class $c$ predicted by model $\mathbf{\omega}^{i}$. $\lambda$ is a hyperparameter balancing the task loss and consistency loss.
The models $\mathbf{\omega}^{i}$ and $\mathbf{\omega}^{j}$ are trained iteratively to minimize their respective losses $\mathcal{L}_{i}$ and $\mathcal{L}_{j}$, leading to mutual enhancement of their predictive performance. This learning paradigm is particularly beneficial in scenarios with model heterogeneity.
4 Methodology
4.1 FedMLAC Overview

### v2

@@ Abstract
Federated Learning (FL) offers a privacy-preserving framework for training audio classification (AC) models across decentralized clients without sharing raw data. However, Federated Audio Classification (FedAC) faces three major challenges: data heterogeneity, model heterogeneity, and data poisoning, which degrade performance in real-world settings. While existing methods often address these issues separately, a unified and robust solution remains underexplored. We propose FedMLAC, a mutual learning-based FL framework that tackles all three challenges simultaneously. Each client maintains a personalized local AC model and a lightweight, globally shared Plug-in model. These models interact via bidirectional knowledge distillation, enabling global knowledge sharing while adapting to local data distributions, thus addressing both data and model heterogeneity. To counter data poisoning, we introduce a Layer-wise Pruning Aggregation (LPA) strategy that filters anomalous Plug-in updates based on parameter deviations during aggregation. Extensive experiments on four diverse audio classification benchmarks, including both speech and non-speech tasks, show that FedMLAC consistently outperforms state-of-the-art baselines in classification accuracy and robustness to noisy data.
@@
[5]organization=Language Technologies Institute, Carnegie Mellon University,city=Pittsburgh, country=USA
[…]
[6]organization=School of Science, Engineering and Environment, University of Salford,city=Manchester, country=UK
@@ 1 Introduction
Audio classification (AC) enables key applications across diverse domains, including speech command recognition [1], emotion detection [2], and environmental sound classification [3]. In practice, however, audio data are often generated and stored locally on edge devices such as smartphones, wearables, and IoT sensors, which raises concerns about data privacy and communication efficiency. These constraints motivate the adoption of Federated Learning (FL), which facilitates collaborative model training without sharing raw data [4, 5], making it a suitable paradigm for real-world audio applications and leading to the emergence of Federated Audio Classification (FedAC) [6]. Despite its potential, FedAC encounters several critical challenges: data heterogeneity resulting from diverse speakers and devices [7], model heterogeneity due to varying client capabilities [8], and data poisoning arising from adversarial inputs [9]. These challenges compromise robustness and scalability, underscoring the need for a unified and resilient learning framework.
@@
While interest in FedAC continues to grow across both speech and non-speech applications, existing research tends to address the challenges of data heterogeneity, model heterogeneity, and data poisoning in partial or disjointed ways. A substantial body of work focuses on mitigating heterogeneous non-IID (non-Independent and Identically Distributed) data distributions through strategies such as regularization techniques [10], client selection [11], representation learning [12], or robust aggregation mechanisms [13, 14]. While these techniques improve global model utility under skewed data, they often assume a homogeneous model architecture shared across all clients, which is impractical in real-world deployments involving heterogeneous edge devices. To address system-level diversity, recent studies have explored flexible model configurations, including personalized layers [15], decoupled training [16, 17], incremental learning [18], or sparsified aggregation [19]. However, such methods rarely account for the interaction between model and data heterogeneity, which can cause unstable convergence, degraded generalization, or fairness concerns when local models diverge significantly.
@@
Another important challenge in FedAC is the presence of corrupted or unreliable local data, commonly referred to as data poisoning. Such corruption can arise not only from adversarial clients but also from natural sources, including label inconsistencies in subjective tasks (e.g., emotion recognition) and distorted audio recordings in uncontrolled environments. These issues can significantly impair model stability and convergence, especially when coupled with non-IID data or heterogeneous client models. To mitigate these effects, researchers have proposed defense strategies such as anomaly detection [20], adversarial learning [21], client clustering [22], and model pruning [23]. While effective in isolation, these methods are typically reactive and rarely consider the joint effects of data and model heterogeneity. Moreover, many introduce a trade-off between robustness and utility, often degrading performance on clean or low-resource clients. Thus, achieving resilience to data corruption without sacrificing personalization or fairness remains an open problem in FedAC.
To address the intertwined challenges of data heterogeneity, model heterogeneity, and data corruption in FedAC, we propose FedMLAC, a Mutual Learning-driven Federated Audio Classification framework designed to achieve both personalization and robustness. FedMLAC comprises two key components: FedMLAC-Update and FedMLAC-Aggregation. To handle data and model heterogeneity, each client maintains a personalized local AC model and a lightweight, globally shared Plug-in model distributed by the server. These models engage in bidirectional mutual learning, where the Plug-in model conveys global knowledge to the local model for generalization, while simultaneously absorbing client-specific information to enhance personalization. This design allows clients with diverse data distributions and model architectures to collaborate effectively without sharing raw data. To mitigate the effects of data corruption caused by adversarial behavior or noisy local inputs, FedMLAC incorporates a robust Layer-wise Pruning Aggregation (LPA) strategy in the server aggregation phase. Instead of naively averaging client updates, LPA selectively prunes anomalous parameter deviations at the layer level, thus filtering out unreliable Plug-in model updates and enhancing the stability and integrity of global aggregation. Together, these two components enable FedMLAC to perform scalable, personalized, and communication-efficient AC learning across heterogeneous FL environments.
The main contributions of this work are summarized as follows:
  • 1.
    We propose FedMLAC, a novel federated audio classification framework that simultaneously tackles the challenges of data heterogeneity, model heterogeneity, and data poisoning, which are rarely addressed in a unified manner by existing approaches.
  • 2.
    We develop a mutual-learning-based local update strategy and a robust layer-wise pruning aggregation (LPA) mechanism that facilitate personalized training while filtering adversarial updates and mitigating client divergence.
  • 3.
    We theoretically analyze the convergence of FedMLAC and empirically demonstrate its effectiveness under heterogeneous and adversarial audio learning conditions.
The remainder of this paper is structured as follows. Section 2 presents a comprehensive review of related work in the field of FedAC. Section 3 describes the proposed approach in detail, including its core components and technical implementation. Section 4 provides a theoretical analysis of convergence. Section 5 outlines the experimental setup and reports the results and analysis across multiple benchmark datasets. Finally, Section 6 concludes the paper by summarizing the key contributions and suggesting directions for future research.
2 Related Work
Recent work in Federated Audio Classification (FedAC) has improved privacy-preserving learning on distributed audio data, but most methods focus on isolated issues. We categorize related efforts around three core challenges: data heterogeneity, model heterogeneity, and data poisoning.
2.1 Addressing Data Heterogeneity
Data heterogeneity is a core challenge in FedAC, as client audio data often differs in content, labels, and recording conditions. Existing works address this through regularization [13, 14], which incorporates proximal terms to reduce divergence between local and global models; client grouping [24, 25], which organizes clients into semantically coherent subgroups; and adaptive training strategies such as dynamic participation scheduling [11] and data-aware model adaptation [10] that improve robustness under non-stationary environments. To enhance generalization under domain shift, [12] proposes contrastive representation learning with weak supervision, while [26] uses local adaptation followed by federated averaging to manage ambient audio diversity. Communication-efficient designs [27, 18] further support heterogeneity by decoupling model components, enabling sparse updates, and relaxing synchronization. While these methods improve FedAC performance under statistical diversity, most still assume consistent model structures and clean inputs, leaving open challenges in addressing both architectural and adversarial variability.
2.2 Handling Model Heterogeneity
Model heterogeneity in FedAC arises from varying client-side capabilities, including compute power, memory, and network bandwidth. To accommodate this, personalized modeling is a key strategy: split architectures [19] allow clients to retain private output heads while sharing a global encoder, and task-specific personalization [15] assigns distinct heads for speaker and emotion classification. To reduce communication and computation costs, methods such as gradient sparsification [28], encoder-head decoupling [29], and bi-level optimization for flexible model configurations [16] have been proposed. Lightweight tuning schemes [30] further adapt acoustic models with minimal overhead. In parallel, asynchronous and adaptive training schedules [18, 27] support flexible client participation and balance communication load in streaming settings. While these approaches enable structural flexibility in FedAC, they often neglect robustness to noisy input conditions.
2.3 Robustness Against Data Poisoning
FedAC is vulnerable to data poisoning, as clients may submit mislabeled, corrupted, or adversarial audio. Existing methods aim to enhance robustness by detecting unreliable updates, filtering noise, or defending against malicious behavior. Anomaly-based defenses play a central role: Tsouvalas et al. [31] use confidence-based outlier detection, while [20] introduces a privacy-preserving verification mechanism via homomorphic encryption and statistical thresholds. Structural defenses modify model behavior to resist compromised updates, such as model pruning with trust scoring [32] and selective gradient masking [23]. Unsupervised methods provide complementary strategies under weak or unlabeled settings. Almaiah et al. [21] utilize signal pre-processing and inter-client validation, while [22] employ federated clustering to detect abnormal label patterns. Earlier work [33] also introduces uncertainty-aware weighting to downscale noisy updates. While these defenses address key vulnerabilities in FedAC, they are rarely integrated with structural personalization or designed for data heterogeneous environments.
2.4 Summary and Motivation
As summarized in Table 1, existing FedAC methods have addressed challenges such as data heterogeneity [24], model diversity [29], and data poisoning [31, 21], but typically in isolation. Few works consider their interaction, despite the fact that in real-world audio applications, these issues often arise together. This fragmentation limits system robustness, fairness, and generalization—highlighting the need for more integrated approaches that can jointly address statistical, structural, and adversarial challenges in FedAC.
@@ 2.4 Summary and Motivation
[13, 14, 24, 12, 26, 10, 11, 25] $\checkmark$ $\times$ $\times$
[28, 19, 15, 16, 30] $\times$ $\checkmark$ $\times$
@@
[34, 29, 27, 18] $\checkmark$ $\checkmark$ $\times$
@@
3 Methodology
3.1 Prelimineries
Problem Setup
@@ 3.1 Prelimineries
where $n=\sum_{k=1}^{N}{n_{k}}$, $n_{k}=|D_{k}|$, $\mathcal{L}_{k}(\omega^{k})=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell(\mathbf{\omega^{k}};\mathbf{x}_{k,j},\mathbf{y}_{k,j})$ denotes the average loss of the ${k^{th}}$ client model $\omega^{k}$ over $D_{k}$, and $\ell(\cdot)$ represents the training loss (e.g., cross-entropy (CE) loss for AC tasks).
@@
$\omega^{k}_{t+(i+1)}=\omega^{k}_{t+(i)}-{\eta_{t+(i)}}\nabla{\mathcal{L}_{k}}(\omega^{k}_{t+(i)}),i=0,1,.,E-1,$ (2)
where $\eta$ is the learning rate, $\omega_{t}^{k}$ represents the $k^{th}$ client model in the $t^{th}$ communication round, and $E$ denotes the local training epochs.
After local training, each client sends its updated model $\mathbf{\omega}_{t+1}^{k}$ to the server, which performs aggregation to update the global model. The most commonly used aggregation method is Federated Averaging (FedAvg) [35], defined as:
@@
The basic FedAC framework requires that each client’s AC model $\mathbf{\omega}^{k}$ maintains the same model architecture. Furthermore, the heterogeneity of distributed AC data leads to an inconsistency between the local learning objectives and the global FedAC training objective. These challenges significantly reduce the practicality and stability of existing FedAC frameworks.
Mutual Learning
@@
$\mathcal{L}_{i}&=\mathcal{L}_{\text{task}}(\mathbf{\omega}^{i};\mathcal{D}_{i})+\lambda\mathcal{L}_{\text{KL}}(\mathit{p}_{i},\mathit{p}_{j})\\ \mathcal{L}_{j}&=\mathcal{L}_{\text{task}}(\mathbf{\omega}^{j};\mathcal{D}_{j})+\lambda\mathcal{L}_{\text{KL}}(\mathit{p}_{j},\mathit{p}_{i})\end{aligned}\right.$ (4)
@@
where $\mathit{p}_{i}(c)$ represents the probability of class $c$ predicted by model $\mathbf{\omega}^{i}$. $\lambda$ is a hyperparameter balancing the task loss and consistency loss. The models $\mathbf{\omega}^{i}$ and $\mathbf{\omega}^{j}$ are trained iteratively to minimize their respective losses $\mathcal{L}_{i}$ and $\mathcal{L}_{j}$, leading to mutual enhancement of their predictive performance. This learning paradigm is particularly beneficial in scenarios with model heterogeneity.
3.2 FedMLAC